OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (0, 6) and a directrix at y = -6.

OpenStudy (anonymous):

@myininaya

myininaya (myininaya):

Hey.

myininaya (myininaya):

so I like to a do a rough drawing to first figure out if my p is going to be negative or positive and where my vertex is going to be.

OpenStudy (anonymous):

it should be positive, just making a mental sketch

myininaya (myininaya):

p is negative if the directrix is above the focus. p is positive if the directrix is below the focus.

myininaya (myininaya):

The vertex is halfway between the vertex and the directrix.

OpenStudy (anonymous):

0,0

myininaya (myininaya):

right

myininaya (myininaya):

and also right on the first thing

myininaya (myininaya):

so we just need to figure out what p is

myininaya (myininaya):

p is the distance between the focus and the vertex

myininaya (myininaya):

or you can look at it as the distance between the vertex and the directrix

OpenStudy (anonymous):

so 6

myininaya (myininaya):

\[4py=x^2 \] is what we had and we knew p was going to be postive because focus was above directrix and yep just replace p with 6

OpenStudy (anonymous):

so 24y=x²

myininaya (myininaya):

yep or you can write it as y=x^2/24 I just divided 24 on both sides

OpenStudy (anonymous):

my answer is y = 1/24x²

myininaya (myininaya):

that sounds good too

OpenStudy (anonymous):

yes, My answer choice is written in the form I wrote it in

OpenStudy (anonymous):

thanks